# Autonomous Stabilization of Floquet States Using Static Dissipation

### v2

@@ Appendix F Stabilization Rate and Fidelity
Both the experimental and theory traces show a clear feature on resonance ($\Delta-B_{0}=0$) with a peak in the steady state $P(\phi_{-})$ accompanied by a steep drop in the time required to reach the steady state. In addition to the single photon resonance between |$\ket{\phi_{+},0}$ and |$\ket{\phi_{-},1}$, both theory and experiment show and additional two-photon resonance at $\delta=-B_{0}/2$, where |$\ket{\phi_{-},2}$ and |$\ket{\phi_{+},0}$ become resonant. This two-photon resonance is a generic feature and persists even outside the adiabatic regime explored in the experiment. Indeed, peaks in the steady state $P(\phi_{-})$ associated with two-photon resonances are visible in Fig. 4 even at high frequency (indicated by dotted lines).
The theory shows good qualitative agreement with the experimental results. Discrepancies in the exact values of the stabilization time and steady-state $P(\phi_{-})$ are likely due to the inability of the Lindblad treatment to capture the colored noise present in our actual experiment. A more complete noise treatment incorporating the experimental noise spectrum is outside the scope of this article.

### v3

@@ Appendix F Stabilization Rate and Fidelity
Both the experimental and theory traces show a clear feature on resonance ($\Delta-B_{0}=0$) with a peak in the steady state $P(\phi_{-})$ accompanied by a steep drop in the time required to reach the steady state, consistent with the minimum stabilization time of roughly $2/\kappa$ predicted by Floquet Lindblad calculation in Appendix D. In addition to the single photon resonance between |$\ket{\phi_{+},0}$ and |$\ket{\phi_{-},1}$, both theory and experiment show and additional two-photon resonance at $\delta=-B_{0}/2$, where |$\ket{\phi_{-},2}$ and |$\ket{\phi_{+},0}$ become resonant. This two-photon resonance is a generic feature and persists even outside the adiabatic regime explored in the experiment. Indeed, peaks in the steady state $P(\phi_{-})$ associated with two-photon resonances are visible in Fig. 4 even at high frequency (indicated by dotted lines).
The theory shows good qualitative agreement with the experimental results. Discrepancies in the exact values of the stabilization time and steady-state $P(\phi_{-})$ are likely due to the inability of the Lindblad treatment to capture the colored noise present in our experiment. In principle, if the noise power spectrum of fluctuations in the qubit frequency is known, alternative simulation techniques such as the quantum trajectories method can be used to exactly model the effect of the colored noise Carmichael (1993). Instead of computing the dynamics of the density matrix, as is done in the Lindblad master equation formalism, these methods compute the exact time dynamics of the system for a particular noise realization, which are then averaged over simulated instances of the noise. The noise spectrum for our device is not sufficiently well-characterized to permit first-principle modeling along these lines. However, our experimental results show that the effect of $T_{2}$ processes are strongly suppressed, to the point that we obtain qualitative agreement using the Lindblad master equation formalism with $T_{1}$ processes alone. In addition, as we showed in related work Ritter et al. (2025), the system decoherence rate $T_{d}$ is independent of the driving field strength $B_{0}$ and rotation rate $\omega_{\mathrm{mod}}$ for a wide range of drive parameters, indicating that the leading order dynamics is well characterized by our simple model.